# Attention is All You Need to Defend Against Indirect Prompt Injection Attacks in LLMs

### v1

@@ I Introduction
In this paper, we propose Rennervate, a defense framework that detects and prevents IPI attacks, as illustrated in Figure 1(b). There are two major challenges facing our design: (1) how to accurately detect stealthy IPI attacks; (2) how to prevent IPI attacks such that the injected instructions are neutralized with little influence on the benign functionality of LLM-integrated applications. On the one hand, IPI attacks are evasive to existing detection methods as the injected instructions may seem benign semantically (e.g., “print Yes”). On the other hand, most existing defense methods cannot perform fine-grained injection sanitization without affecting benign instructions. Furthermore, most existing IPI defense methods fail to transfer to unseen attacks. Therefore, it is crucial to design defense methods that achieve robust IPI detection based on inherent features and fine-grained IPI sanitization to maintain the functionalities of LLM-integrated applications.
To address these challenges, we adopt attention features to detect and prevent covert IPI attacks. Attention features have shown great potential in analyzing LLM behaviors [yu2024unveiling, zhao2024explainability, zhang2023tell]. By examining the characteristic attention distributions of the target LLM associated with instruction processing, Rennervate identifies the act of task injection regardless of the specific purpose of injected tasks, which forms a more robust and transferable IPI detection and sanitization framework. Attention features vary in length due to different prompt/response token lengths, making it challenging to extract stable information. To tackle this difficulty, we introduce a token-level detector, which enables fine-grained IPI detection. To enhance the transferability of Rennervate, we propose a 2-step attentive pooling mechanism, which aggregates attention heads and response tokens based on their importance for injection analysis. We parallelize all token-level detectors to achieve lightweight detection. Furthermore, we construct a large-scale IPI dataset with fine-grained token-level labels featuring various IPI attacks and NLP tasks. The dataset, named FIPI, will be open-sourced to boost research on IPI attack defenses.
@@ 15    $InjLst\leftarrow\emptyset$;
Algorithm 1 Detection and Localization.

### v2

@@ I Introduction
In this paper, we propose Rennervate, a defense framework that detects and prevents IPI attacks, as illustrated in Figure 1(b). The design of Rennervate faces two primary challenges: (1) how to accurately detect semantically stealthy IPI attacks; (2) how to effectively neutralize injected instructions while preserving the benign functionality of LLM-integrated applications. Firstly, because injected instructions can appear semantically benign (e.g., “Please print Yes.”), they often evade conventional detection methods. Secondly, most existing defenses lack the fine-grained capability to sanitize malicious injections without impacting benign instructions. Furthermore, current methods frequently fail to generalize to unseen attack variants. Therefore, it is crucial to design defense methods that achieve robust IPI detection based on inherent features and fine-grained IPI sanitization to maintain the functionalities of LLM-integrated applications.
To address these challenges, we leverage attention features for the detection and prevention of IPI attacks. Attention mechanisms have demonstrated considerable potential in interpreting and diagnosing LLM behaviors [yu2024unveiling, zhao2024explainability, zhang2023tell]. By analyzing the distinctive attention patterns that a target LLM exhibits when processing instructions, Rennervate can identify task-injection attempts, thereby establishing a more robust and transferable framework for IPI detection and sanitization. However, since attention features vary in prompt/response length, extracting stable and generalizable information is non-trivial. To overcome this, we introduce a token-level detector that enables fine-grained analysis of potential injections. To further improve the transferability of Rennervate, we design a 2-step attentive pooling mechanism that aggregates information across attention heads and response tokens according to their relevance to injection analysis. All token-level detectors are parallelized to ensure lightweight and efficient detection. Additionally, we construct a large-scale IPI dataset, named FIPI, which contains fine-grained token-level annotations across a diverse set of IPI attacks and NLP tasks. This dataset will be made publicly available to support further research in IPI defense.
@@ 15    $InjLst\leftarrow\emptyset$;
Algorithm 1 Detection and Sanitization.